Left Direct Gregorian Technical Observing with IRTC: UT 2009 October 12

Observer: JHill (LBTO), DMiller (Tucson), JEsguerra (LBTO) plus JStorm(Potsdam)
Telescope Operator: Aaron Ceranski
Instrument Support: TShih
Telescope Support: BCurtis
SW support: (Tucson)

Summary

Had troubles with GCS/AZCAM/AGW4 at the beginning. Lost 20 minutes. Don't know what was wrong.

Got 16-17 pointing stars starting in evening twilight (but fit to the Oct 9 data was better).

Closed for a cap cloud at 03:15 UT. Reopen at 11:00 UT.

More problems with GCS/AZCAM/AGW4 when we reopen with just the same symptoms as the beginning of the night. Lost 30 minutes.

The bad news is that the "new" transformation is giving poor results with 5 arcsec errors.

This week has been so bad, that John has now changed his opinion - we no longer have too much LDG commissioning time on the schedule.

Details

All Times in UT

00:45 pointto StoneLagw gs=0 ACQUIRE with dome closed. But GCS fails to get the acquire image.

00:50 Same result on second try.

00:53 Stop and restart GCSL

00:54 Acquire preset succeeds

Open

00:58 Open T=4.2 degC D=-2.1 degC Wind gusting to 10 m/sec. A few puffy clouds scattered about.

01:01 pointto M5_1539 gs=0 TelM="ACQUIRE" PAM- "ERROR: Guide or movie image didn't arrive within t...." (as above)

01:04 Stop and restart GCSL (as above)

01:05 pointto M5_1539 gs=0 TelM="ACQUIRE" PAM- couldn't move probe to position (SFP) 1.41059e-08, 0.117918

01:06 Stop GCSL, stopAGW -u 4 (RPC error)

01:08 sudo reboot (lbtmu105)

01:11 startAGW succeeds immediately (without waiting 5-10 minutes as last night). So then stopAGW, startAGW to cycle the power as we originally started out to do

01:13 home the motors for AGW4. (RPC communication error, but then it proceeded to home ??)

01:16 start GCS

01:17 pointto M5_1539 gs=0 TelM="ACQUIRE" PAM- Found star at IE=-8 CA=43

01:20 pointto M5_1539 gs=2 TelM="ACTIVE" PA=INDEF

01:24 Clear the Secondary Offload that we forgot to do at the start. This leaves us with a big defocus. Initialize lookup tables, clear residual primary Active Optics. (and the guider followed all that)

01:30 Clear secondary Active Optics as it seems to be diverging.

01:38 Finally converged to 0.7 arcsec images.

01:39 pointto M5_1539 gs=0 TelM="ACQUIRE" PAM- - Adjust IE=-12 CA=31

Pointing Data

(We recognize that the very start of the night is not the best time to collect stable pointing data, but this is the last night of this run, and we are running short of sky hours.)

Data in /Repository/IRTC_Data/091012

Start pointing log PTLog20091011IRTCdirect. IRTC is using J filter.

01:45 gopoint 1 active+ ploop+

(The 20 sec hardwired delay in the gopoint script is just a bit too short to allow for active optics corrections. We are using the guide acquisition to pull in stars from the edge of the field.)

01:51 missed the star in grid 2 irtc.20091012.015207.fits cube152 is empty.

02:00 missed the star in grid 5 irtc.20091012.020139.fits cube155 is empty.

02:14 missed the star in grid 8 irtc.20091012.021437.fits cube158 is empty. (There was a glint in the acq image.)

02:22 missed the star in grid 11 irtc.20091012.022211.fits cube161 is empty. (We are missing the stars below elevation 25 deg.)

02:24 star in grid 15 wasn't obvious on IRTC, but was there on guider. Clouds? Or we got the wrong star?

02:38 Stop after grid 17 as the star is getting out of focus.

02:39 pointto M5_1539 gs=2 TelM="ACTIVE" PA=INDEF - IE and CA are within 1 arcsec of where we set them 1 hour ago. Seeing is 0.75 arcsec.

02:42 takeseq 0.1 100 frm=10 on M5_1539 so this can be a pointing datum also. irtc.20091012.024259.fits

02:44 gopoint 2 active- ploop- / ptspiral IE=-12 CA=61 / takeseq irtc.20091012.024842.fits

02:49 gopoint 5 ACT0281 / takeseq irtc.20091012.025202.fits

02:52 gopoint 8 WT10_068 / takeseq irtc.20091012.025717.fits cube 171

02:58 gopoint 11 WT10_225 irtc.20091012.030006.fits cube172

03:01 pointto M5_0850 gs=0 TelM="ACQUIRE" PAM- Adjust IE=-20 CA=58

03:04 pointto StoneOagw gs=11 TelM="ACTIVE" PA=0 - No star found. But there was a faint shadow. Clouds? Yes, it has become "soupy".

03:07 pointto StoneOagw gs=0 TelM="ACQUIRE" PAM- - Star is there, but then fades away.

Results Made a pointing model with 13 stars with RMS=9.09. Copied to /home/telescope/TCS/Configuration/PCS/irtcdg.20091012.ptmod

       coeff       change     value    sigma

  1     IA       -274.944   -274.94   41.934
  2     IE         -0.882     -0.88    6.950
  3     NPAE       -9.350     -9.35   46.486
  4     CA        +24.413    +24.41   58.034
  5     AW         -4.007     -4.01    3.116
  6     AN        +21.809    +21.81    3.015
  7     TF        +67.184    +67.18    9.697
  8     HACA2      -5.846     -5.85    4.581

Sky RMS =   9.09
Popn SD =  14.66

  • tpoint Graphs:
    irtcdg.20091012.ptmod.jpg

Close

03:17 Close for a cap cloud forming over the mountain. T=3.3 degC D=-0.5

Stopped pointing log.

03:25 Direct Rotator has hit a relative limit and tripped the MCR. MCSPU thinks the rotator is at 184:28:30 which doesn't seem to be near a limit. Aaron and Brian are working on it. Called Mike G. This is a relative limit between the rotator and the cable chain. Aaron held down the bypass switch, closed MCR, and we were able to do a "rcready ldg".

05:06 Humidity has reached 100%

Open

11:00 Humidity has dropped rapidly so we reopen to clear sky. T=1.9 D=-14.2 !!!

11:10 pointto M5_0256 gs=0 ACQUIRE - "ERROR: Guide or movie image didn't arrive within t...."

11:12 Stop and restart GCS

11:13 pointto M5_0256 gs=0 ACQUIRE -

11:15 GCS died while we were adjusting IE/CA.

11:16 pointto M5_0256 gs=0 ACQUIRE - Adjust IE=-16 CA=48

11:18 pointto StoneCagw gs=18 TelM="ACTIVE" PA=0 - didn't find guide star.

11:21 pointto StoneCagw gs=0 ACQUIRE PAM-

Install new pointing model

11:24 stop and restart PCS with new pointing model irtcdg.20091009.ptmod that Doug analyzed during the "fog period".

11:24 pointto M5_0151 gs=0 ACQUIRE - GCS doesn't get the image again

Oct 12 11:24:56 lbtmu304 LBT_left_GCS: PROFILING: requesting image...
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: AzCam: calculateTimeout4Image: Region: 1, 1, 560, 528
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: AzCam: calculating timeout for image from PEPSI1 with size of 593280 bytes and 2000 ms exposureTime to be 4748 ms
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: IST: btr = 1
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: possibly damaged FITS file /tmp/guide.fits: END tag not found
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: CamImage::receiveGuideImage() : error in header of FITS file
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: IST: error when receiving image
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: IST: closing open socket
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: IST: socket closed
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: IST: status = -1
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: AzCam: IST reported error while receiving image.
Oct 12 11:24:56 lbtmu304 LBT_left_GCS: error: MovieThread: acquisition image didn't arrive within timeout of 6748 ms.

11:25 stop and restart GCS

11:26 pointto M5_0151 gs=0 ACQUIRE - Preset failed because can't move probe. Aarghh.

Oct 12 11:25:52 lbtmu304 LBT_left_GCS: AGWUnitAIP::doSetProbePosition: calling oacontrol::setxy(agw-control, 3, x:-0.692138 y:612.262854, 23.000000, 0) ...
Oct 12 11:26:53 lbtmu304 LBT_left_GCS: AGWUnit: timeout occured waiting for probe to reach destination. pthread_kill result is 0
Oct 12 11:26:53 lbtmu304 LBT_left_GCS: AGWUnit: cancelling worker thread: -1452565616
Oct 12 11:26:53 lbtmu304 LBT_left_GCS: AGWUnit: cancellation succeeded.
Oct 12 11:26:53 lbtmu304 LBT_left_GCS: gcs.interface.agw.setProbePosition.failed: couldn't move probe to position (SFP) -0.105059, 0.153209, focus: 33.3876 :timed out waiting for probe to reach destination
Oct 12 11:26:53 lbtmu304 LBT_left_GCS: error: setProbePosition: couldn't move probe to position (SFP) (x,y,z): -0.105059, 0.153209, 33.387599 mm.
Oct 12 11:26:53 lbtmu304 LBT_left_GCS: gcs.command.moveProbeOnGuideStar.failed: AGW move failed
Oct 12 11:26:53 lbtmu304 LBT_left_GCS: gcs.command.setGuidestarList.failed: guidestar list successfully changed but probe couldn't be moved to position

11:27 Stop GCS. stopAGW gives RPC error. Reboot lbtmu105.

11:30 stopAGW gives RPC error now.

11:33 startAGW (localhost: RPC: program not registered) same as yesterday.

11:35 (after waiting 5 minutes) stopAGW / startAGW (wait 2 minutes) home motors

11:40 start GCSL. Note that GCSGUI always reports "starting up" but never reports "ready"

11:41 pointto M5_0151 gs=0 ACQUIRE - Adjust IE=0 CA=72

11:47 pointto StoneCagw gs=0 ACQUIRE PAM- hit the star dead nuts.

11:48 pointto StoneCagw gs=18 TelM="ACTIVE" PA=0 - star is 2 sec from the hotspot. Seeing is 0.9 to 1.3 arcsec at el 43 deg. Need 2 sec exposure on IRTC to see it.

11:53 takeseq 2 20 irtc.20091012.115401.fits cube173 (oops, forgot to do a getdat on that one - it is repeated in the next script.)

11:55 cl < home$tran_StoneC_00.cl
transtar (10) irtc.20091012.115704.fits cube174
transtar (14)
transtar (15)
transtar (12)
transtar (18)

We can see that this "new" transform is worse than the one we replaced. Called Jesper. He is rechecking the configuration file.

12:06 cl < home$tran_StoneC_90.cl
transtar (21) irtc.20091012.120807.fits cube179
transtar (20)

12:11 cl < home$tran_StoneC_180.cl
transtar (13) - didn't find star at all irtc.20091012.121301.fits is empty
transtar (16) ditto
Abort the script

12:16 pointto StoneCagw gs=18 ACTIVE found 8 arcsec to left.

12:17 cl < home$tran_StoneC_270.cl
transtar (9) star way out in corner of acq irtc.20091012.121911.fits cube182
transtar (10) irtc.20091012.122116.fits cube183
transtar (3) star was off acq, but guider got it irtc.20091012.122326.fits cube184
Aborted script.

12:24 pointto StoneCagw gs=0 ACQUIRE PAM- Adjust IE=-7 CA=62

12:27 cl < home$tran_StoneC_315.cl
IRAF pipe error during pointto
transtar (9) irtc.20091012.122809.fits cube185

Script dies, so restart it.

12:29 cl < home$tran_StoneC_315.cl
transtar (9) irtc.20091012.122943.fits cube186
Script crashes.

Start new IRAF window. Lost 2 minutes.

12:32 cl < tran_StoneC_190.cl
first star not found Aborted

We can't find stars near PA=180, but in StoneC these are all at high radius in the field. So maybe we have some kind of angle problem.

12:34 cl < home$tran_StoneC_315.cl
transtar (9) irtc.20091012.123618.fits meteor in the IRTC cube187
transtar (10) irtc.20091012.123823.fits cube188
transtar (3) irtc.20091012.124030.fits cube189
transtar (14) irtc.20091012.124243.fits cube190
transtar (15)
transtar (12) star wasn't found

Dice5 data

This is pushing into twilight so we pick the brightest guidestar in StoneC near the edge of the field.

12:47 pointto StoneCagw gs=25 ACTIVE

12:49 dice5 7 7 DETXY 2 30 loop=2 This starts with irtc.20091012.124947.fits and cube192.
Second step fails because probe couldn't move that far. The star really is at the edge of the field. Fifth step fails also so we only have half of the die.

12:58 WFS is too bright for the second loop through the dice, so image quality is poor. Loop 2 position 1 was irtc.20091012.125732.fit Aborted.

All the data including stacked cubes lives in /Repository/IRTC_Data/091012_Transform

Close

13:00 Close dome for impending dawn. T=2.8 degC D=-16.8 degC Wind 4 m/sec. Sky is clear with some pretty crepuscular rays in the east.

From PCSInstrument.conf
INSTRUMENT              IRTCDG  
FOCALSTATION            directGregorian
LEFTPOINTINGMODEL       irtcdg.20091009.ptmod
LEFTFOCALLENGTH         123421.4
LEFTPOINTINGORIGIN      0.0 0.0
LEFTXYFLIP              true false
LEFTIAA                 0.0
LEFTPAI                 0.0
LEFTWAVELENGTH          1.400
LEFTZEROPOINT           +4.45
LEFTSCALE               -1.0

-- DougMiller - 11 Oct 2009
I Attachment Action Size Date Who Comment
daq.20091012.gifgif daq.20091012.gif manage 46 K 14 Oct 2009 - 01:22 JohnHill Plot of Telescope and Mirror Temperatures
irtcdg.20091012.ptmod.jpgjpg irtcdg.20091012.ptmod.jpg manage 41 K 12 Oct 2009 - 08:15 DougMiller tpoint Graphs
Topic revision: r7 - 08 Jan 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback